OpenStudy (anonymous):

Differentiate. f(x)=3x(2x+3)^5

OpenStudy (anonymous):

You have to use both chain and product rules. \[f'(x)=3x \times 5 \times (2x+3)^4 \times 2 + (2x+3)^5 \times 3\]and simplify...

OpenStudy (anonymous):

so f'(x)= 3(2x+3)^4 [10x+2x+3] f'(x)=3(2x+3)^4 [12x+3] f'(x)=9(2x+3)^4 [4x+1]

OpenStudy (anonymous):

Yep.